OpenStudy (anonymous):

I need help with hyperbolas please? Posting the hyperbola equations and questions in the comments.

OpenStudy (anonymous):

x^2 - y^2 = 1

OpenStudy (anonymous):

\[\frac{ (x-1)^{2} }{ 2 }-\frac{ (y+3)^{2} }{ 2 }=1\]

OpenStudy (anonymous):

The center of the hyperbola is (1, -3), right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Also, I need to find the lengths of the transverse and conjugate axes, the slopes of the asymptotes, and the foci, but I'm really not quite sure how to do that...

OpenStudy (anonymous):

the slope of the asymptotes is the easiest part if you have \[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\] the slopes are \(m=\pm\frac{a}{b}\) in this case \(\pm1\)

OpenStudy (anonymous):

Oh my, I wrote the equation wrong, a^2 should be 64 and b^2 should be 81.

OpenStudy (anonymous):

So, then the slopes of the asymptotes will be 64/81? or would that be vice versa, 81/64? is a always greater than b?

OpenStudy (anonymous):

ok no problem in this case you have \[\frac{(x-1)^2}{8^2}-\frac{(y+3)^2}{9^2}=1\]

OpenStudy (anonymous):

the slope is not \(\pm\frac{a^2}{b^2}\) it is \(\pm\frac{a}{b}\)

OpenStudy (anonymous):

in your case \(\pm\frac{8}{9}\)

OpenStudy (anonymous):

does that need to be simplified any further or can it stay as a fraction?

OpenStudy (anonymous):

actually it is wrong

OpenStudy (anonymous):

should be \(\pm\frac{9}{8}\) i had it upside down

OpenStudy (anonymous):

oh, okey dokey. so does that need to be simplified or can it stay as a fraction?

OpenStudy (anonymous):

not sure what you mean by "simplify" there is no common factors, so you cannot reduce it just leave it as it is

OpenStudy (anonymous):

I mean't, like, turn it into a decimal, but I figured it should just be left as is, so thank you (:

OpenStudy (anonymous):

Is there a formula I use to find the foci and the lengths of the axes?

OpenStudy (anonymous):

i think the lengths are just 8 and 9

OpenStudy (anonymous):

that is, semi major is 8 semi minor is 9 not much to that part

OpenStudy (anonymous):

finding the foci starts with having an idea of what this looks like

OpenStudy (anonymous):



OpenStudy (anonymous):

there is my lousy picture, but i drew it to show that the foci are to the right and left of the center, which is \((1,-3)\)

OpenStudy (anonymous):

Yes, I know where they're located, I'm just not sure how to find their coordinates using the equation.

OpenStudy (anonymous):

since \(a^2+b^2=64+81=145\) you know \(c=\sqrt{145}\)

OpenStudy (anonymous):

and since the center is \((1,-3)\) the foci are at \((1-\sqrt{145},-3)\) and \((1+\sqrt{145},-3)\)

OpenStudy (anonymous):

Oh, okay, so the square root of 145 is 12.04159458, so that means the foci are (-11.04159458, -3) and (13.04159458, -3), right?

OpenStudy (anonymous):

Is it okay to round those numbers? Will it still be accurate enough?

OpenStudy (anonymous):

wait a second, is this hyperbola vertical? i remember reading something a while back saying that a (the major axis) is always longer (greater) than b (the minor axis), and if they're flipped, wouldn't that make it vertical?

OpenStudy (anonymous):

yeah i drew it wrong answer was correct though you can check it here http://www.wolframalpha.com/input/?i= \frac{%28x-1%29^2}{8^2}-\frac{%28y%2B3%29^2}{9^2}%3D1

OpenStudy (anonymous):

Okay! Thank you!